Two small nonconducting spheres have a total charge of 94.0 μC . When placed 1.11 m apart, the force each exerts on the other is 11.5 N and is repulsive. What is the charge on each?

What if the force were attractive?
Answer #1

a) for repulsive force

q1 + q2 = 94

F = k q1 q2 / r^2

11.5 = 9*10^9*10^-12* ( 94 - q2) q2 / 1.11^2

-q2^2 + 94 q2 - 1574.35 = 0

solving for q2

q2 = 21.8 uC

q1 = 72.2 uC

===========

for attractive

q1 - q2 = 94

11.5 = 9*10^9*10^-12 * ( 94 + q2) q2 / 1.11^2

q2^2 + 94 q2 - 1574.35 = 0

q2 = 14.51 uC

q1 = 108.51 uC

=========

Comment in case any doubt, will reply for sure. Goodluck